OpenStudy (anonymous):

Silva set up a dog-walking business. He earned $60 in June. He is projected to earn 110% more in July. How much is Silva projected to make in July? $126 $170 $176 $660

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

Please help me!!!!!!!!!

OpenStudy (michele_laino):

I think that gain of Silva, in July, is: \[60+\frac{ 110 }{ 100 }*60=60+\frac{ 11 }{ 10 }*60=60+66=126$\] so A.

OpenStudy (anonymous):

thx, could you help me with some more??

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

At a water park, a one-day ticket costs $72.50 per adult. Which equation can be used to calculate the one-day ticket price, W, for n adults? n = 72.50W W equals 72.50 over n n equals 72.50 over W W = 72.50n

OpenStudy (michele_laino):

I think that price for W=72.50*n because for one adult price is W=72.50*1, for two adults, price is: W=72.50*2 for three adults, price is W=72.50*3, and so on....

OpenStudy (michele_laino):

yes it is!

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

ill open another question

OpenStudy (michele_laino):

ok!